Mary'age is ⅔ that of Peter's. Two years ago Mary's age was ½ of what Pete 's age will be in 5 years'time. How old is Peter now?
Answer by Boreal(15235)

Peter is now 27.
Peter's age is x
Mary's is (2/3)x
[(2/3)x]-2= (1/2)(x+5)
distribute
(2/3)x-(2)=(1/2)x +(5/2)
Multiply everything by 6
4x-12=3x=15
x=27, Peter's age
Mary is 18
Two years ago, she was 16 and that is half of Peter's age in 5 years, or 32.